OpenStudy (anonymous):

John is packing books into boxes. Each box can hold either 15 small books or 8 large books. He needs to pack at least 35 boxes and at least 350 books. Write a system of linear inequalities to represent the situation.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

show me what you have

OpenStudy (anonymous):

I wish I knew how to solve this, but I don't. Please help.

OpenStudy (anonymous):

x+y>=35 x+y>=350 I think?

jimthompson5910 (jim_thompson5910):

hmm let me think for a moment

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

15x+8y>=35 15x+8y>=350?

jimthompson5910 (jim_thompson5910):

I'm thinking maybe x = number of boxes y = number of books `He needs to pack at least 35 boxes and at least 350 books` so, x >= 35 y >= 350 I don't know how to deal with the `Each box can hold either 15 small books or 8 large books` information

OpenStudy (anonymous):

Ok, thank you!

jimthompson5910 (jim_thompson5910):

no problem